Treasury Stock Transactions

Lava Lake Inc. bottles and distributes spring water. On February 11 of the current year, Lava Lake reacquired 6,000 shares of its common stock at $72 per share. On April 30, Lava Lake Inc. sold 4,200 of the reacquired shares at $78 per share. On August 22, Lava Lake Inc. sold 1,800 shares at $68 per share.

a. Journalize the transactions of February 11, April 30, and August 22. For a compound transaction, if an amount box does not require an entry, leave it blank.

Feb. 11
Apr. 30
Aug. 22

b. What is the balance in Paid-In Capital from Sale of Treasury Stock on December 31 of the current year?
$  

c. For what reasons might Lava Lake have purchased the treasury stock?
Answer #1
a
Feb. 11 Treasury Stock 432000 =6000*72
       Cash 432000
Apr. 30 Cash 327600 =4200*78
     Treasury Stock 302400 =4200*72
     Paid-In Capital from Sale of Treasury Stock 25200
Aug 22 Cash 122400 =1800*68
Paid-In Capital from Sale of Treasury Stock 7200
     Treasury Stock 129600 =1800*72
b
Paid-In Capital from Sale of Treasury Stock balance 18000 Credit =25200-7200
c
For reissuance to employees as a bonus according to stock purchase agreements.
To support the market price of the stock
To provide shares for resale to employees